OpenStudy (anonymous):

6. How many real number solutions does the equation have? y = 2x² – 20x + 50 one solution two solutions no solutions infinitely many solutions

OpenStudy (solomonzelman):

set it equal to zero (instead of y) and apply the quadratic formula.

OpenStudy (anonymous):

I don't get your formula

OpenStudy (solomonzelman):

or actually not a quadratic, factor it out after setting it equal to zero.

OpenStudy (solomonzelman):

\[\huge\color{blue}{ 2x^2-20x+50=0 }\]\[\huge\color{blue}{ 2(x^2-10x+25)=0 }\]\[\huge\color{blue}{ (x^2-10x+25)=0 }\] factor

OpenStudy (anonymous):

your making it compilacting

OpenStudy (solomonzelman):

Look at my work in blue.

OpenStudy (anonymous):

still stuck

OpenStudy (solomonzelman):

can you factor out,\[\huge\color{green}{ x^2-10x+25= 0 }\] ( it's a perfect square )

OpenStudy (solomonzelman):

what 2 numbers add to get 10, and multiply to 25 ?

OpenStudy (anonymous):

Do you just want an answer?

OpenStudy (shamil98):

it's probably from a quiz, since it is multiple choice.... anyways GOFORMITTENS 2k14

OpenStudy (anonymous):

Because you don't want to answer any of @SolomonZelman 's questions.

OpenStudy (anonymous):

There is no solution because those terms can't add up.

OpenStudy (anonymous):

\[y=2\left( x ^{2} -10x+25\right)=2\left( x-5 \right)^{2}\ge 0 \]